Bicycling, the worlds leading cycling magazine, reviews hundreds of bicycles throughout the year. Their Road-Race category

Bicycling, the world's leading cycling magazine, reviews hundreds of bicycles throughout the year. Their "Road-Race" category contains reviews of bikes used by riders primarily interested in racing. One of the most important factors in selecting a bike for racing is the weight of the bike. The following data show the weight (pounds) and price ($) for 10 racing bikes reviewed by the magazine.

(a) Use the data to develop an estimated regression equation that could be used to estimate the price for a bike given the weight. (Round your numerical values to the nearest integer). 

(b) Computer r2. (Round your answer to three decimal places.) 

(c) Predict the price (in dollars) for a bike that weighs 16 pounds. (Round your answer to the nearest dollar.)
Answer #1

Answer:

a) y = 28574 - 1439 (x)

b) r^2 = 0.864

The estimated regression equation provides a good fit, since r^2 > or = 0.55

c) Price for a bike that weights 16 pounds = 28574 - 1439 (16) = $5550